# ACC-Debate: An Actor-Critic Approach to Multi-Agent Debate

### v1

@@ Abstract
Large language models (LLMs) have demonstrated a remarkable ability to serve as general-purpose tools for various language-based tasks. Recent works have demonstrated that the efficacy of such models can be improved through iterative dialog between multiple models, frequently referred to as multi-agent debate (MAD). While debate shows promise as a means of improving model efficacy, most works in this area treat debate as an emergent behavior, rather than a learned behavior. In doing so, current debate frameworks rely on collaborative behaviors to have been sufficiently trained into off-the-shelf models. To address this limitation, we propose ACC-Debate, an Actor-Critic based learning framework to produce a two-agent team specialized in debate. We demonstrate that ACC-Debate  outperforms SotA debate techniques on a wide array of benchmarks.
@@ 1 Introduction
In particular, multi-agent debate approaches make use of off-the-shelf general-purpose LLMs, which are not trained to collaborate. Such approaches rely on collaboration as an emergent, rather than a learned, behavior. While, in some cases, these emergent behaviors are sufficient, the question remains: Can these methods be improved by imbuing models directly with collaborative abilities? To answer this, we propose a novel paradigm to train teams of LLMs to collaboratively solve tasks.
Most relevant to the main idea of this paper is DebateGPT (Subramaniam et al., 2024). DebateGPT uses debate as a mechanism to attain higher-quality fine-tuning data. Differing from our work, their approach focuses on using debate to generate better training data for a single model that gives single responses, rather than optimizing the LLMs themselves for collaborative problem-solving over multiple rounds of conversation.
In this paper, we propose a novel framework Actor-Critic Debate (ACC-Debate) which jointly trains a two-agent team to collaboratively solve problems through iterative conversation; this team consists of an actor-agent, responsible for providing answers for a given task, and a critic-agent, responsible for assisting the actor-agent with feedback on its answers. In our training pipeline, we introduce a novel off-policy learning scheme called ”guided-debate” to generate high-quality multi-turn training data to enhance the actor’s and critic’s performance on challenging tasks.
@@
  • We are the first to propose a framework for training a team of LLMs jointly (Actor-Critic) within the context of debates.
  • We introduce a novel data generation scheme, “guided debate trajectories”, which enables the efficient creation of high-quality training data for both the actor and critic roles.
  • Our extensive experiments demonstrate that our method, ACC-Debate, significantly outperforms existing state-of-the-art approaches.
@@ 2 Related Work
Our research is closely related to the emerging field of multi-agent debate (sometimes called multi-LLM deliberation) which examines how to use groups of models to solve tasks through iterative discussion Chan et al. (2023); Liang et al. (2023); Du et al. (2023); Khan et al. (2024); Michael et al. (2023); Rasal (2024); Pham et al. (2023); Abdelnabi et al. (2023); Hong et al. (2023); Irving et al. (2018); Li et al. (2023b; c; 2024a); Wang et al. (2023a); Zhang et al. (2023). Many of these works find that language models have naturally collaborative abilities Singhal et al. (2023); Du et al. (2023); Chan et al. (2023), while others have noted that the collaborative ability of off-the-shelf models can be quite limited Wang et al. (2024); Smit et al. .
@@
Current approaches to multi-agent debate can be broadly cast into two main categories: those that modify model prompts and responses during the debate Liang et al. (2023); Khan et al. (2024); Rasal (2024); Feng et al. (2024); Yang et al. (2024), and those that modify the structure of the debate process Li et al. (2023a); Hong et al. (2023); Liu et al. (2023); Li et al. (2024c); Wang et al. (2023b); Wu et al. (2023); Chen et al. (2023); Chang (2024b). Importantly, both categories use off-the-shelf language models (which have not been trained to collaborate) and work by modifying either the inputs or outputs of these models. Deviating from this line of work, we aim to specifically train a team of models to collaboratively solve tasks.
@@
Two works of particular note are that of Subramaniam et al. (2024), which proposes to use debate data to fine-tune models, and Li et al. (2024b), which trains models for adversarial debate. In the former, debate is used to generate higher-quality fine-tuning data and is not used at inference time; differing from this work, we train models directly to collaborate and use debate both during training and inference. In the latter, models are trained to be effective arguers rather than collaborators, i.e., models are trained to give conceiving arguments such that they can win a debate against other LLMs. Differing from this work, we train models to collaboratively solve tasks.
@@
In the context of multi-agent deliberation, the concept of divergent opinions is highly relevant to our method. Several approaches to multi-agent debate aim to control the level of disagreement among the agents Liang et al. (2023); Khan et al. (2024); Chang (2024a). Often, these works dynamically increase disagreement to prevent early convergence of debate. In our study, we leverage divergent opinions to generate high-quality training data. In particular, we have debaters change their opinion during the debate and measure whether or not that change increases or decreases the likelihood that the debate converges to a correct answer; this is used to asses the value of a given training example.
@@
Also closely related to our work are paradigms that aim to use self-generated data to improve model performance, often in the context of reasoning or chain of thought Trung et al. (2024); Huang et al. (2024); Xiong et al. (2024); Chen et al. (2024); Pang et al. (2024b); Zelikman et al. (2022). Similar to this line of research, we make use of model generations as training data. However, we are the first work to use such data in the context of multiple models debating collaboratively to solve a task.
@@ 3 Preliminaries and notation
In this section, we formalize the debate between an Actor (a model that provides answers) and a Critic (a model that provides feedback to the actor) while also introducing the notation that will be used throughout the remainder of the paper.
Let $(x,y)\sim\mathcal{D}$ be a task-answer pair source from a distribution of tasks and answers $\mathcal{D}$. For a given task $x$, two models – an actor model responsible for providing answers and a critic model responsible for providing feedback and assistance to the actor model – engage in an iterative discussion over $T$ rounds, to correctly infer the answer $y$. Let $\theta_{a}$ and $\theta_{c}$ be the parameters of actor and critic models, respectively. The iterative discussion between these two models is as follows:
@@
The accuracy of this procedure is measured via the correctness of the actor’s final response, i.e., $\mathbb{I}\big{[}\zeta(z_{a}^{(T)})=y\big{]}$. Where $\zeta$ is a function that extracts answers from text-based responses. For example if $z_{a}^{(T)}=$ “The sky is blue”, then $\zeta(z_{a}^{(T)})=$“blue”. With this notation and formalization of debate, we introduce our framework for training actor-critic teams to debate.
Figure 1: ACC-Debate training pipeline.
@@ 4 Methodology
In this section, we outline our procedure for training a two-agent debate team, consisting of an actor model $f_{\theta_{a}}$ (responsible for providing answers to a given task $x$) and a critic model $f_{\theta_{c}}$ (responsible for providing feedback and assistance to the actor). At inference time, the two trained agents engage in iterative discussion to solve a given task $x$, generating the final response $z_{a}^{(T)}$.
4.1 An Actor-Critic Debate Framework
@@
$\theta_{a}^{*},\,\theta_{c}^{*}=\arg\max_{\theta_{a}}\max_{\theta_{c}}\mathbb{E}_{(x,y)\sim D}\bigg{[}\zeta\bigg{(}\underbrace{f_{\theta_{a}}\big{(}x,z_{a}^{(T-1)},z_{c}^{(T-1)}\big{)}}_{\text{actor's final response}z_{a}^{(T)}}\bigg{)}=y\bigg{]}$ (1)
Intuitively, Eq. 1 aims to simultaneously optimize the actor’s parameters $\theta_{a}$ and the critic’s parameters $\theta_{c}$, ensuring that the actor’s final output at iteration $T$ matches the correct answer $y$. In other words, we optimize the accuracy of the actor’s response at time $T$, namely $z_{a}^{(T)}=f_{\theta_{a}}\left(x,z_{a}^{(T-1)},z_{c}^{(T-1)}\right)$, where accuracy is measured as $\mathbb{E}\big{[}\zeta\big{(}z_{a}^{(T)}\big{)}=y\big{]}$.
It is important to note that the recursive nature of multi-agent debate introduces significant complexity to the optimization process. Each response $z_{a}^{(t)}$ depends not only on the actor’s previous output $z_{a}^{(t-1)}$ but also on the critic’s previous output $z_{c}^{(t-1)}$. This interaction closely resembles a cooperative dynamic Stackelberg game (Li & Sethi, 2017), where two players engage in hierarchical decision-making over time, leading us to adopt an iterative best-response approach (Fiez et al., 2019). In other words, we first train the critic model, followed by training the actor to best respond to the critic’s output. We can then update the critic to adapt to the newly trained actor, and so on. More formally, this process works by first fixing $\theta_{a}$, and solving,
@@
this process then repeats until a desired stopping criteria is reached. In practice, we find that a single iteration is sufficient to produce a high-quality debate team.
While this alternating scheme allows us to optimize the actor and critic separately, the objectives of each model still cannot be optimized directly due to the recursive nature of model responses in this objective; responses at round $T$ depend on those given by the model at round $t-1$ which themselves depend on the response given at round $t-2$ and so on. To deal with this temporal dependency, we next introduce the concept of Partial Trajectory rewards, which will allow us to capture the signal of each response $z^{(t)}$ for each $t\leq T$.
@@ 4.2 Partial Trajectory Reward
To address the inter-round dependencies of the above optimization, we proposed a scheme that allows us to determine the “goodness” of a given response $z^{(t)}$ (from either the actor or the critic) for any $t\leq T$. Consider a conversation between the actor and the critic that was paused at time $t$, i.e., the most recent response is $z^{(t)}$. To assess the goodness of $z^{(t)}$, one might ask how likely the debate procedure will converge to the correct answer $y$ at round $T$, given that the procedure is already at response $z^{(t)}$. Formally, we can define this as
$r(z^{(t)},x,y)=\mathbb{E}\bigg{[}\zeta(z_{a}^{(T)})=y|~{}x,z^{(t)}\bigg{]}$
Intuitively, the partial reward captures the expectation of arriving at the correct answer $y$ through debate starting at round $t$ with generation $z^{(t)}$. In practice, $r(z^{(t)},x,y)$ can be estimated by learning the reward structure $r$ or by using heuristics such as one-step roll-out, i.e., Monte Carlo estimation. Empirically, we observe that one-step roll-out heuristics are effective in producing high-quality training data and have the added benefit of being significantly more efficient than learning-based approaches.
Our objective will then be to optimize the actor and critic, $\theta_{a},\theta_{c}$, so that the responses produced by these models at each timestep $t$, namely $z^{(t)}$, maximize $r(z^{(t)},x,y)$. That is, we optimize the actor and the critic so that at each timestep $t$, they give a response $z^{(t)}$ which has a high probability of leading the debate to converge to the correct answer at time $T$.
@@
Data: Actor and critic: $\theta_{a},\theta_{c}$, Distribution of tasks $\mathcal{D}$
@@
       $\mathbf{z}^{(0)}\leftarrow$ OneDebateRound$(x)$   /* actor and critic responses, i.e. $\mathbf{z}=\langle z_{a}^{(0)},z_{c}^{(0)}\rangle$/
@@
             $\mathbf{z}^{(t)}\leftarrow$ OneDebateRound$(x,\mathbf{z}^{(t-1)})$   /* updated natural responses*/
            /* guided-debate towards, and away from, correct answer $y$/
             $\mathbf{z}_{+}^{(t)}\leftarrow$ OneGuidedDebateRound$(x,\mathbf{z}^{(t-1)},y)$
@@
            /* Estimate final round accuracy if debate continue from response $z^{(t)}$,
@@
            /* Compute the expected improvement in debate for each trajectory */
             /* Save trajectory pairs that result in sufficient accuracy improvment */
             if $v_{+}-v\geq t$ then
@@ 4.3 Off-Policy Trajectory Generation
In this section, we describe how to generate the preference data needed to optimize the objective in Eq. 1. The classification of a sample as positive or negative is determined by the debate trajectory it follows. Specifically, a positive sample for training the actor corresponds to a trajectory likely to lead to the correct answer at round $T$, while a negative sample corresponds to one that leads to an incorrect answer at round $T$. Intuitively, we aim to push the actor model to generate responses that lead to correct answers while reducing responses that are unlikely to do so, thus optimizing for Eq. 3. The same principle applies to the critic when optimizing Eq. 2.
@@
With this intuition in mind, we now describe how such data is generated. A debate trajectory can be defined as a sequence of responses $\langle z_{a}^{(0)},z_{c}^{(0)},\ldots,z_{a}^{(T)},z_{c}^{(T)}\rangle$ for a given task $x$. One way to generate preference data is to generate multiple rollouts at each round and select the trajectories with the highest (lowest) value of $r(z^{(t)},x,y)$ as positive (negative) samples. This approach could enforce the desired behavior for both the actor and the critic if enough samples are collected.
@@
However, this approach is not without its limitations. In particular, if the model performs poorly on a given dataset, it may be difficult to collect enough positive samples, resulting in low training signals. Additionally, even if the model performs adequately, generating sufficient responses for both the actor and critic requires significant computational resources, especially to ensure that high $r(z^{(t)},x,y)$ values are used for positive samples and low values for negative samples.

### v3

# ACC-Collab: An Actor-Critic Approach to Multi-Agent LLM Collaboration

@@ Abstract
Large language models (LLMs) have demonstrated a remarkable ability to serve as general-purpose tools for various language-based tasks. Recent works have demonstrated that the efficacy of such models can be improved through iterative dialog between multiple models. While these paradigms show promise in improving model efficacy, most works in this area treat collaboration as an emergent behavior, rather than a learned behavior. In doing so, current multi-agent frameworks rely on collaborative behaviors to have been sufficiently trained into off-the-shelf models. To address this limitation, we propose ACC-Collab, an Actor-Critic based learning framework to produce a two-agent team (an actor-agent and a critic-agent) specialized in collaboration. We demonstrate that ACC-Collab  outperforms SotA multi-agent techniques on a wide array of benchmarks.
@@ 1 Introduction
footnotetext: Code available at https://github.com/LlenRotse/ACC-Collab
@@
Similar to most multi-agent paradigms, MAD approaches make use of off-the-shelf general-purpose LLMs, which are not trained to collaborate. Such approaches rely on collaboration as an emergent, rather than a learned, behavior. While, in some cases, these emergent behaviors are sufficient, the question remains: Can these methods be improved by imbuing models directly with collaborative abilities? To answer this, we propose training teams of LLMs to solve tasks collaboratively.
A particularly relevant work is DebateGPT (Subramaniam et al., 2024), which employs debate as a mechanism to generate higher-quality fine-tuning data. Unlike our approach, which optimizes LLMs for multi-round collaborative problem-solving, their method focuses on using debate to enhance training data for a single model that produces individual responses.
In this paper, we propose a novel framework Actor-Critic Collaboration (ACC-Collab) which jointly trains a two-agent team to collaboratively solve problems through iterative conversation; this team consists of an actor-agent, responsible for providing answers for a given task, and a critic-agent, responsible for assisting the actor-agent with feedback on its answers. In our training pipeline, we introduce a novel off-policy learning scheme called ”Guided-Collaboration” to generate high-quality multi-turn training data to enhance the actor’s and critic’s performance on challenging tasks.
@@
  • We are the first to propose a framework for jointly training a team of LLM agents (Actor-Critic) within the context of collaborative problem solving.
  • We introduce a novel data generation scheme, “Guided Collaboration Trajectories”, which enables the efficient creation of high-quality training data for both the actor and critic roles.
  • Our extensive experiments demonstrate that our method, ACC-Collab, significantly outperforms existing state-of-the-art approaches.
@@ 2 Related Work
Our research is closely related to the emerging field of multi-agent deliberation, sometimes called Multi-Agent Debate (MAD), which examines how to use groups of models to solve tasks through iterative discussion Chan et al. (2023); Liang et al. (2023); Du et al. (2023); Li et al. (2023c); Khan et al. (2024); Michael et al. (2023); Rasal (2024); Pham et al. (2023); Abdelnabi et al. (2023); Hong et al. (2023); Irving et al. (2018); Li et al. (2023b; d; 2024a); Wang et al. (2023a); Zhang et al. (2023). Many of these works find that language models have naturally collaborative abilities Singhal et al. (2023); Du et al. (2023); Chan et al. (2023), while others have noted that the collaborative ability of off-the-shelf models can be quite limited Wang et al. (2024); Smit et al. .
@@
Current approaches to multi-agent deliberation can be broadly cast into two main categories: those that modify model prompts and responses during the discussion Liang et al. (2023); Khan et al. (2024); Rasal (2024); Feng et al. (2024); Yang et al. (2024), and those that modify the structure of the deliberation process Li et al. (2023a); Hong et al. (2023); Liu et al. (2023); Li et al. (2024c); Wang et al. (2023b); Wu et al. (2023); Chen et al. (2023); Chang (2024b). Importantly, both categories use off-the-shelf language models (which have not been trained to collaborate) and work by modifying either the inputs or outputs of these models. Deviating from this line of work, we aim to specifically train a team of models to collaboratively solve tasks.
@@
Two works of particular note are that of Subramaniam et al. (2024), which proposes to use debate data to fine-tune models, and Li et al. (2024b), which trains models for adversarial debate. In the former, debate is used to generate higher-quality fine-tuning data and is not used at inference time; differing from this work, we train models directly to collaborate and use multi-agent discussion both during training and inference. In the latter, models are trained to be effective arguers rather than collaborators, i.e., models are trained to give conceiving arguments such that they can win a debate against other LLMs. Differing from this work, we train models to collaboratively solve tasks.
@@
In the context of multi-agent deliberation, the concept of divergent opinions is highly relevant to our method. Several approaches to multi-agent deliberation aim to control the level of disagreement among the agents Liang et al. (2023); Khan et al. (2024); Chang (2024a). Often, these works dynamically increase disagreement to prevent early convergence of deliberation. In our study, we leverage divergent opinions to generate high-quality training data. In particular, we have agents change their opinion during the discussion and measure whether or not that change increases or decreases the likelihood that the agents’ discussion converges to a correct answer. Using this signal we can then asses the value of a given training example for training the models.
@@
Also closely related to our work are paradigms that aim to use self-generated data to improve model performance, often in the context of reasoning or chain of thought Trung et al. (2024); Huang et al. (2024); Xiong et al. (2024); Chen et al. (2024); Pang et al. (2024b). Similar to this line of research, we make use of model generations as training data. However, we are the first work to use such data in the context of multiple models debating collaboratively to solve a given task.
@@ 3 Preliminaries and notation
In this section, we formalize multi-agent collaboration between an Actor (an agent that provides answers) and a Critic (an agent that provides feedback to the actor) while also introducing the notation that will be used throughout the remainder of the paper.
Let $(x,y)\sim\mathcal{D}$ be a task-answer pair source from a distribution of tasks and answers $\mathcal{D}$. For a given task $x$, two agents – an actor agent responsible for providing answers and a critic agent responsible for providing feedback and assistance to the actor agent – engage in an iterative discussion over $T$ rounds, to correctly infer the answer $y$. Let $\theta_{a}$ and $\theta_{c}$ be the parameters of actor and critic agent, respectively. The iterative discussion between these two agents is as follows:
@@
The accuracy of this procedure is measured via the correctness of the actor’s final response, i.e., $\mathbb{I}\big{[}\zeta(z_{a}^{(T)})=y\big{]}$. Where $\zeta$ is a function that extracts answers from text-based responses. For example if $z_{a}^{(T)}=$ “The sky is blue”, then $\zeta(z_{a}^{(T)})=$“blue”. With this notation and formalization of multi-agent collaboration, we introduce our framework for training actor-critic teams.
Figure 1: ACC-Collab training pipeline, exemplified for the actor. 1) We generate data from both natural deliberation as well as guided deliberation towards and away from the ground truth answer $y$ using the actor and critic. 2) We compute the relative quality of each trajectory based on the expected quality difference $\Delta_{y},\Delta_{!y}$ w.r.t. to the natural response. 3) We store all high-quality pairwise data in our database and train the actor agent. 4) We alternate this procedure for the actor and critic. See Figure 5 of the supplement for the corresponding procedure applied to the critic.
@@ 4 Methodology
In this section, we outline our procedure for training a two-agent team, consisting of an actor agent $f_{\theta_{a}}$ (responsible for providing answers to a given task $x$) and a critic agent $f_{\theta_{c}}$ (responsible for providing feedback and assistance to the actor). At inference time, the two trained agents engage in iterative discussion to solve a given task $x$, generating the final response $z_{a}^{(T)}$.
4.1 An Actor-Critic Collaboration Framework
@@
$\theta_{a}^{*},\,\theta_{c}^{*}=\arg\max_{\theta_{a}}\max_{\theta_{c}}\mathbb{E}_{(x,y)\sim D}\bigg{[}\zeta\bigg{(}\underbrace{f_{\theta_{a}}\big{(}x,z_{a}^{(T-1)},z_{c}^{(T-1)}\big{)}}_{\text{actor's final response}~{}z_{a}^{(T)}}\bigg{)}=y\bigg{]}$ (1)
Intuitively, Eq. 1 aims to simultaneously optimize the actor’s parameters $\theta_{a}$ and the critic’s parameters $\theta_{c}$, ensuring that the actor’s final output at iteration $T$ matches the correct answer $y$. In other words, we optimize the accuracy of the actor’s response at time $T$, namely
$z_{a}^{(T)}=f_{\theta_{a}}\left(x,z_{a}^{(T-1)},z_{c}^{(T-1)}\right),$
where accuracy is measured as $\mathbb{E}\big{[}\zeta\big{(}z_{a}^{(T)}\big{)}=y\big{]}$.
It is important to note that the recursive nature of multi-agent deliberation introduces significant complexity to the optimization process. Each response $z_{a}^{(t)}$ depends not only on the actor’s previous output $z_{a}^{(t-1)}$ but also on the critic’s previous output $z_{c}^{(t-1)}$. This interaction closely resembles a cooperative dynamic Stackelberg game (Li & Sethi, 2017), where two players engage in hierarchical decision-making over time, leading us to adopt an iterative best-response approach (Fiez et al., 2019). In other words, we first train the critic agent, followed by training the actor to best respond to the critic’s output. We can then update the critic to adapt to the newly trained actor, and so on. More formally, this process works by first fixing $\theta_{a}$, and solving,
@@
this process then repeats until a desired stopping criteria is reached. In practice, we find that a single iteration is sufficient to produce a high-quality collaborative team.
While this alternating scheme allows us to optimize the actor and critic separately, the objectives of each agent still cannot be optimized directly due to the recursive nature of agent responses in this objective; responses at round $T$ depend on those given by the agent at round $t-1$ which themselves depend on the response given at round $t-2$ and so on. To deal with this temporal dependency, we next introduce the concept of Partial Trajectory rewards, which will allow us to capture the signal of each response $z^{(t)}$ for each $t\leq T$.
@@ 4.2 Partial Trajectory Reward
To address the inter-round dependencies of the above optimization, we proposed a scheme that allows us to determine the “goodness” of a given response $z^{(t)}$ (from either the actor or the critic) for any $t\leq T$. Consider a conversation between the actor and the critic that was paused at time $t$, i.e., the most recent response is $z^{(t)}$. To assess the goodness of $z^{(t)}$, one might ask how likely the deliberation procedure will converge to the correct answer $y$ at round $T$, given that the procedure is already at response $z^{(t)}$. Formally, we can define this as
$r(z^{(t)},x,y)=\mathbb{E}\bigg{[}\zeta(z_{a}^{(T)})=y|~{}x,z^{(t)}\bigg{]}$ (4)
Intuitively, the partial reward captures the expectation of arriving at the correct answer $y$ through deliberation starting at round $t$ with generation $z^{(t)}$. In practice, $r(z^{(t)},x,y)$ can be estimated by learning the reward $r$ or by using heuristics such as one-step roll-out, i.e., Monte Carlo estimation.
In our experiments, we use one-step roll-out heuristics, i.e. simulating an additional deliberation round multiple times from response $z^{(t)}$. The reward $r(z^{(t)},x,y)$ is set as the average accuracy of these simulations. Empirically, we find this approach effective for generating high-quality training data. We leave learning-based reward functions for future work.
Our objective will then be to optimize the parameters of the actor and critic, $\theta_{a},\theta_{c}$, so that the responses produced by these agents at each timestep $t$, namely $z^{(t)}$, maximize $r(z^{(t)},x,y)$. That is, we optimize the actor and the critic so that at each timestep $t$, they give a response $z^{(t)}$ which has a high probability of leading the deliberation to converge to the correct answer at time $T$.
@@
Data: Actor and critic: $\theta_{a},\theta_{c}$, Distribution of tasks $\mathcal{D}$ , Reward threshold $\varepsilon$
@@
       $\mathbf{z}^{(0)}\leftarrow$ OneDeliberationRound$(x)$   /* actor and critic responses, i.e. $\mathbf{z}=\langle z_{a}^{(0)},z_{c}^{(0)}\rangle$/
@@
             $\mathbf{z}^{(t)}\leftarrow$ OneDeliberationRound$(x,\mathbf{z}^{(t-1)})$   /* updated natural responses*/
            /* guided-deliberation towards, and away from, correct answer $y$/
             $\mathbf{z}_{+}^{(t)}\leftarrow$ OneGuidedDeliberationRound$(x,\mathbf{z}^{(t-1)},y)$
@@
            /* Estimate final round accuracy if deliberation continue from response $z^{(t)}$,
@@
            /* Compute the expected improvement for each trajectory */
             /* Save trajectory pairs that result in sufficient accuracy improvement */
             if $v_{+}-v\geq\varepsilon$ then
@@ 4.3 Off-Policy Trajectory Generation
In this section, we describe how to generate the preference data needed to optimize the objective in Eq. 1. The classification of a sample as positive or negative is determined by the deliberation trajectory it follows. Specifically, a positive sample for training the actor corresponds to a trajectory likely to lead to the correct answer at round $T$, while a negative sample corresponds to one that leads to an incorrect answer at round $T$. Intuitively, we aim to push the actor agent to generate responses that lead to correct answers while reducing responses that are unlikely to do so, thus optimizing for Eq. 3. The same principle applies to the critic when optimizing Eq. 2.
@@
With this intuition in mind, we now describe how such data is generated. A deliberation trajectory can be defined as a sequence of responses $\langle z_{a}^{(0)},z_{c}^{(0)},z_{a}^{(1)},z_{c}^{(1)},\ldots,z_{a}^{(T)},z_{c}^{(T)}\rangle$ for a given task $x$. A straightforward way to generate preference data would be to generate multiple rollouts at each round and select the trajectories with the highest $r(z^{(t)},x,y)$ as positive samples and those with the lowest $r(z^{(t)},x,y)$ as negative samples. This approach could enforce the desired behavior for both the actor and the critic if enough samples are collected.
@@
However, this approach is not without its limitations. In particular, if the agent performs poorly on a given dataset, it may be difficult to collect enough positive samples, resulting in low training signals. Additionally, even if the agent performs adequately, generating sufficient responses for both the actor and critic requires significant computational resources, especially to ensure that high $r(z^{(t)},x,y)$ values are used for positive samples and low values for negative samples.
4.4 Guided Collaborative Trajectories